6) The length of a rectangle is three times the breadth. If the minimum perimeter of the rectangle is 160 cm, then what can you say about breadth? a) Breadth = 20 b) Breadth 20 c) Breadth 20 d) Breadth 20

6)    The length of a rectangle is three times the breadth. If the minimum perimeter of the rectangle is 160 cm, then what can you say about breadth?
a)    Breadth = 20
b)    Breadth  20
c)    Breadth  20
d)    Breadth  20

Grade:upto college level

2 Answers

V. Harshit -
10 years ago

length=3x    breadth=x

2(l+b)=2(3x+x)=8xwhich is the perimeter

8x=160cm

x=20cm

so, breadth=20 cm

length=60 cm

sri harsha
10 years ago

a) breadth = 20